OpenStudy (lovelycharm):

solve each system using elimination 3x - 2y =1 8x + 3y = 2

OpenStudy (anonymous):

YOU CAN MULTIPLY THE FIRST EQAUTION BY3 AND THE SECOND EQAUTION BY 2

OpenStudy (anonymous):

3x - 2y =1 8x + 3y = 2 (6)(3x-2y = 1) (4)(8x+3y = 2) 18x - 12y = 6 32x +12y = 8 40x = 14 x = 14/40 = 7/20 plug x into either equation to find y

OpenStudy (anonymous):



OpenStudy (anonymous):

DO YOU GET WHAT IM DOING?

OpenStudy (lovelycharm):

no

OpenStudy (anonymous):

inorder to eliminate you need to have the opposite number

OpenStudy (lovelycharm):

i got it so x = 7 over 25

OpenStudy (anonymous):

okay now plug that in

OpenStudy (lovelycharm):

how i found y

OpenStudy (anonymous):

your right its not 7/20 its my answer well if you go to freemathhelp.com theyll give u the answer!

OpenStudy (anonymous):

whats y?

OpenStudy (anonymous):

y should be -2/25

OpenStudy (lovelycharm):

u have to find y and x i find x how u find y

OpenStudy (anonymous):

u subsitute inx to find y

OpenStudy (lovelycharm):

u could gave me the step u found that answer

OpenStudy (anonymous):

kk

OpenStudy (anonymous):